CURRICULUM VITAE

Richard Wiegandt was born on September 19, 1932 in Budapest, Hungary. His father was the manager of a linen factory who died in 1939. After the grammer school he studied at the University of Szeged and graduated in 1955 as a high school teacher in mathematics and physics. He started with scientific research in algebra as a student of Professor L. Rédei.

After graduation he was assigned to teach in the remote town Orosháza, and worked there as a high school teacher from 1955 till 1961. His application for a Ph.D. fellowship was turned down five times, in 1959 – when he had already published three research papers in English – also the reason was disclosed: he had been found politically not suitable for scientific work. After having published two more research papers in German, in 1961 he was accepted for Ph.D. studies (aspiranture) at the L. Eötvös University in Budapest under the supervision of Professor L. Fuchs.

From 1964 till 2007 he worked at the Mathematical Institute of the Hungarian Academy of Sciences as a scientific co-worker, scientific chief co-worker (1970), scientific adviser (1978) and between 1981 and 1996 he acted as head of Department of Algebra. He tought also compulsory and optional courses at the L. Eötvös University (1962-1975) and at the Technical University Budapest (1975-1978).

Scientific qualifications:

1 Dr. rer. nat. conferred by the L. Eötvös University, Budapest, 1967; 2 Candidate of Math. Sci. conferred by the Hungarian Academy of Sciences,

1967; 3 Doctor of Math. Sci. conferred by the Hungarian Acedemy of Sciences,

1975.

Longer visits at foreign institutions:

1. Mathematical Institute of the Polish Academy of Sciences Warsaw, Ph.D.

study leave, 6 months, 1964/65. He was visiting professor at several universities on all the five continents.

2. University of Islamabad, Pakistan, UNESCO-visiting professor (specialist in algebra, grade P-5, the highest in the professional category) from 1970 upto 1972.

Subsequent visiting professorships:

3. Technische Universität Wien, Austria, 1 month 1974;

4. Science Centre at the University of Alexandria, Egypt, 6 months 1976/77;

5. Technische Universität Clausthal, FR Germany, Wintersemester 1978/79;

6. University of British Columbia Vancouver, Canada, 2 months 1979 ;

7. University of Tasmania Hobart, Australia, 2 months 1980;

8. Cairo University, Egypt, 1 month 1982; 9. University of British Coulumbia Vancouver, Canada, fall semester, 1982

(teaching calculus);

10. Dalhousie University Halifax, Canada, 2 months 1983;

11. University of Pretoria, South Africa, 2 months 1983;

12. University of Florida Gainesville, USA, spring semester 1986 (teaching differential equations);

13. University of Tasmania Hobart, Australia, 1 month 1987;

14. University of the Orange Free State Bloemfontein, South Africa, 3 months 1990;

15. National Cheng Kung University Tainan appointed by the National Science Council at Taipei, Taiwan, R.O.C., 12 months 1991/92;

16. University of Southwestern Louisiana Lafayette, USA, spring semester 1995 (teaching survey of calculus and decision mathematics);

17. Johannes Kepler Universität Linz, Austria, Wintersemester 1997/98;

18. Universität Bielefeld and Technische Universität Clausthal, FR Germany, 1 month 1998.

He applied for and got a Humboldt-Fellowship for 1 year in Germany 1968, but for political reasons he was not allowed to accept it. Upon this basis he got the Humboldt-Sonderforschungsstipendia:

19. Universität Düsseldorf, TU Clausthal and Universität Tübingen, 3 months 1987;

20. TU Clausthal, 1 month 1991;

21. TU Clausthal, 1 month 1993;

22. TU Clausthal, 1 month 2000.

Conferences and lectures:

He participated on many international conferences and gave talks at several universities in the following countries: Austria, Australia, Brazil, Bulgaria, Canada, China, Czechoslovakia (Czechia, Slovakia), East-Germany, Egypt, Estonia, Germany, Holland, Hong Kong, Hungary, Iran, Israel, Italy, Japan, Moldova, Mongolia, Pakistan, Poland, Portugal, Romania, Slovakia, Slovenia, South-Africa, Soviet Union (Estonia, Latvia, Moldova, Russia, Ukraine), Spain, Ukraine, United Kingdom (England, Scotland), United States, Taiwan, Thailand, Turkey, Vietnam, West-Berlin, West-Germany, Yugoslavia (Bosnia-Hercegovina, Serbia).

Research:

His main research field has been algebra (rings, radicals, categories, semigroups, near-rings, automata), but published papers also in combinatorics, graph theory and topology.

Scientific publications are listed separately.

Coauthors:

Ottó Steinfeld (Hungarian), Ferenc A. Szász, (Hungarian), Eqbal Ahmed (Bengali), Muneer A. Rashid (Pakistani), Alexandr V. Arhangel’skiĭ (Russian), Ervin Fried (Hungarian), Mario Petrich (American), László Márki (Hungarian), Patrick N. Stewart (Canadian), Johann Wiesenbauer (Austrian), Leo C. A. van Leeuwen (Dutch), Cornelis Roos (Dutch), Alfred Widiger (German), László Rédei (Hungarian), William G. Leavitt (American), Chehata G. Chehata (Egyptian), Pham Ngoc Ánh (Vietnamese), Tim Anderson (Canadian), Juriĭ M. Rjabuhin (Russian-Moldavian), Barry J. Gardner (Australian), Gerhard Betsch (German), Wilfried Lex (German), Nguyen Van Loi (Vietnamese), Ismail A. Amin (Egyptian), Adam Suliński (Polish), Mike M. Parmenter (Canadian), Ben de la Rosa (South African), Stefan Veldsman (South African), Kalle Kaarli (Estonian), Rainer Mlitz (Austrian), Andor Kertész (Hungarian), Konstantin I. Beĭdar (Russian), Julia S. van Niekerk (South African), Hanns Joachim Weinert (German), Fong Yuen (Taiwanese), Lee Pjek-hwee (Taiwanese), Huang Feng-kuo (Taiwanese), Arthur D. Sands (British), Sodnomkhorloo Tumurbat (Mongolian), Gary F. Birkenmeier (American), Henry E. Heatherly (American), Enoch K. S. Lee (Hong Kong), Usama A. Aburawash (Egyptian), Liu Shao-xue (Chinese), Edmund R. Puczyłowski (Polish), D. Isabel C. Mendes (Portuguese), Bhavanari Satyanarayana (Indian), Jan Krempa (Polish).

Supervisor:

1 Dinh Van Huynh, Ph.D. 1975, Matrin–Luther–Universität, Halle GDR; 2 M. Sadiq Zia, dr. rer. nat. 1975, L. Eötvös University, Budapest; 3 Pham Ngoc Ánh, C.Sc. 1982, Hungarian Academy of Sciences, Budapest; 4 Tran Lam Hách, C.Sc. 1984, Hungarian Academy of Sciences, Budapest; 5 Nguyen Van Loi, C.Sc. 1988, Hungarian Academy of Sciences, Budapest; 6 Usama A. Aburawash, C.Sc. 1991, Hungarian Academy of Sciences, Bu-

dapest; 7 Gábor Hetyei, Ph.D. 1994, MIT, USA; 8 Sodnomkhorloo Tumurbat, Ph.D. 1993, Ulaan Baatar, Mongolia; 9 Mátyás Domokos, C.Sc. 1996, Hungarian Academy of Sciences, Budapest.

Member of editorial board of journals:

1. Studia Scientiarum Mathematicarum Hungarica, 1981-1996;

2. Karachi Journal of Mathematics, 1984-1989;

3. Pure Mathematics and Applications, (Budapest), 1990-1994;

4. Quaestiones Mathematicae (South Africa), 1991-2002;

5. Beiträge zur Algebra und Geometrie (Berlin, Heldermann–Verlag), 1991-2002;

6. Journal of Egyptian Mathematical Society, 1993-2002;

7. Algebra Colloquium (Beijing and Singapore, Springer–Verlag), 1994-2002;

8. Mathematica Pannonica (Leoben, Pécs and Trieste), 1998-2002;

9. Mathematical Notes, University of Miskolc (Hungary), 2000-2002.

Reviewer of the Zentralblatt für Mathematik (1965-2002) and of the Mathematical Reviews (1970-2002).

Languages: fluent in English and German, reads Russian, native language Hungarian.

Membership: J. Bolyai Math. Soc. since 1955, American Math. Soc. since 1984; South African Math. Soc. (honorary member) since 1990.

He has never been a member of any political party.

Married since 1957, has two sons, Peter (1960) and Thomas (1966).

Budapest, January 1, 2012.